what number decreased by 33 1/3% is 30?
16 is 25% more than what number?
An item sells for $45 after a 20% discount. Find the original price.
What is 66 2/3% less than 45?
A price plus 6% tax is equal to $4.24. Find the original price.

Answer by stanbon(75887)

what number decreased by 33 1/3% is 30?
Let "x" be the number:
x-(33 1/3)x =30
0.667x= 30
x = 44.98
--------------------
16 is 25% more than what number?
Let the number be "x".
16 = 0.25x
x = 64
------------
An item sells for $45 after a 20% discount. Find the original price.
Let the original price be "x"
x - (20%)x = 45
0.80x = 45
x=?
I'll leave that to you.
------------------------
What is 66 2/3% less than 45?
Let "x" be the answer
x = 45 - (0.667)x
(1+0.667)x = 45
(1.667)x = 45
x = ?
---------------------
A price plus 6% tax is equal to $4.24. Find the original price.
Let original price be x.
x + 0.06x=4.24
1.06x = 4.24
x = ?
===============
Cheers,
Stan H.